OpenStudy (anonymous):

1) Write an equation of a line in slope intercept form with slope -3 going through the point (2, - 2). Show your work for full credit.

OpenStudy (anonymous):

They give you a point and the slope. So, use Point-Slope form. Then change to slope-intercept form. Point-Slope Form: y - y1 = m(x - x1) y - (-2) = -3(x - 2) y + 2 = -3(x - 2) Change to slope intercept form: y + 2 = -3(x - 2) y + 2 = -3x + 6 y = -3x + 4 Hope this helps! (:

OpenStudy (anonymous):

@timothyford